# A Selective Homomorphic Encryption Approach for Faster Privacy-Preserving Federated Learning

### v2

@@ IV-F1 Homomorphic Encryption
In the context of federated learning, HE enables secure aggregation of model updates from distributed clients. The mathematical foundation of HE ensures that:
@@
where Enc and Dec denote encryption and decryption, respectively, and Eval represents a computation on encrypted data. This property is critical for privacy-preserving applications, as it prevents adversaries from accessing plaintext data [16].
IV-F2 Differential Privacy
Differential privacy (DP) is a privacy-preserving framework that introduces controlled noise to statistical computations, ensuring that the presence or absence of a single data point does not significantly influence the output.
Formally, a mechanism $M$ satisfies $\epsilon$-differential privacy if for any two adjacent datasets $D_{1}$ and $D_{2}$ differing in a single entry, and for any subset of outputs $O$:
@@ IV-F2 Differential Privacy
where $\epsilon$ is the privacy budget, controlling the trade-off between privacy and utility [30].
In federated learning, DP is used to add noise to model updates or gradients before aggregation, protecting individual contributions while preserving the overall utility of the model. This makes DP a valuable tool for maintaining privacy in sensitive domains such as healthcare and finance.
IV-G Process Flow of the Proposed Model
The proposed model operates in three main stages: client-side processing, server-side processing, and client post-processing. Each stage has a set of steps, illustrated in pseudocode and explained in detail below.
IV-H Federated Learning System Process Flow
Figure 2 illustrates the client-side process, while Figure 3 shows the server-side process in the federated learning system with encryption and scrambling techniques.
Figure 1: General depiction of the federated learning model showing the processes of encryption, scrambling, and aggregation.
Figure 1 provides a general depiction of our federated learning model, demonstrating the encryption, scrambling, and noise addition steps across client nodes, and the subsequent aggregation performed by the server.
Figure 2: Client-side process in federated learning system: encryption, scrambling, and noise addition.
Figure 3: Server-side process in federated learning system: aggregation and re-scrambling.
The process begins at the client side, where local data is used to train a model, producing a weight vector. This weight vector is split, encrypted, and further obfuscated through bitwise scrambling and the addition of noise.
Once encrypted, the weight vectors are formatted to hexadecimal and transmitted to the server. The server then receives encrypted data from all clients and performs federated averaging (FedAvg) to aggregate the model parameters. Importantly, the aggregation is done on both encrypted and non-encrypted vectors, ensuring privacy is maintained while also allowing for more efficient computation.
@@ IV-H Federated Learning System Process Flow
This secure and privacy-preserving process enables effective model training across decentralized data while preventing unauthorized access to sensitive information.
IV-H1 Client-Side Processing
In the client-side processing stage, model weights ($w$) are prepared for secure transmission to the server by encrypting selected weights, obfuscating the remaining weights, and combining them into a single data structure. Initially, the weights are split into two subsets based on a predefined encryption percentage ($enc\_pct$) (Lines 6–8). The selected portion of the weights is encrypted using a homomorphic encryption function ($enc\_func$), and these encrypted weights are stored in $enc\_w$ (Lines 9–11). For the remaining weights, they are first formatted to resemble encrypted data ($fmt\_weight$), noise is added using $noise\_func$, and the noisy weights are scrambled with a cryptographic key ($scr\_key$) via a scrambling function ($scr\_func$) (Lines 12–16). These scrambled weights are stored in $scr\_w$. Finally, the encrypted and scrambled weights are combined into a single structure ($enc\_scr\_w$) (Line 17) and transmitted to the server (Line 18).
@@ IV-H1 Client-Side Processing
IV-H2 Server-Side Processing
@@
IV-H3 Client Post-Processing
In the final stage, the client receives the processed weights ($sec\_w$) from the server (Line 2) and restores them for updating the local model. The encrypted subset of weights is decrypted using the appropriate decryption function (Lines 3–4), while the scrambled weights are unscrambled with the cryptographic key ($scr\_key$) to restore their original form (Lines 5–7). These two subsets of weights are then combined to reconstruct the full model weights (Line 8), which are subsequently used to update the client’s local model for the next round of federated learning (Line 9).
@@ IV-H3 Client Post-Processing
15:     Proceed to next round of federated learning with $final\_w$
@@
IV-I Advantages of the Proposed Model
@@
  • Scalability: Optimized for large-scale federated learning systems with multiple clients.
IV-J Privacy Analysis
In this section, we formally show that our federated learning mechanism, which combines selective homomorphic encryption, differential privacy, and bitwise scrambling, maintains a well-defined differential privacy guarantee.
IV-J1 Setup and Definitions
We consider a federated learning setting with $n$ clients, each holding private data. The global model is described by parameters indexed by $[N]$. Let $S\subseteq[N]$ be the subset of parameters to be protected by selective homomorphic encryption (HE), and let $[N]\setminus S$ be the remaining parameters to which we add differential privacy (DP) noise.
@@ IV-J1 Setup and Definitions
A mechanism $M$ is $\epsilon$-DP if for any two adjacent datasets $D,D^{\prime}$ (differing in one record) and any event $O$:
@@
Parameters in $S$ are encrypted using a semantically secure HE scheme. Semantic security ensures no polynomial-time adversary can distinguish ciphertexts of different messages, implying no additional privacy cost in a DP sense (0-DP).[16]
@@
For each $i\in[N]\setminus S$, we add noise calibrated to $\epsilon_{i}$-DP. By the composition property of DP, releasing all these parameters together is $\left(\sum_{i\in[N]\setminus S}\epsilon_{i}\right)$-DP.[30]
@@
We define a scrambling function $\mathcal{T}_{k}:\mathcal{X}\to\mathcal{X}$, where $\mathcal{X}$ is the space of possible model outputs. $\mathcal{T}_{k}$ is deterministic, keyed by $k$, and independent of $D$ except through $M(D)$. This scrambling is a form of post-processing.[43]
IV-J2 Privacy Guarantee
Theorem 1.
@@ IV-J2 Privacy Guarantee
    Encrypts $W_{i}$ for $i\in S$ using a semantically secure HE scheme [16, 44].
@@
    Adds noise to each $W_{j}$ for $j\in[N]\setminus S$ to achieve $\epsilon_{j}$-DP. Releasing all noisy parameters together is $\left(\sum_{j\in[N]\setminus S}\epsilon_{j}\right)$-DP [42].
@@
    Applies a scrambling function $\mathcal{T}_{k}$ to the entire output, resulting in $\mathcal{M}^{\prime}(D)=\mathcal{T}_{k}(\mathcal{M}(D))$ [43].
@@
Proof.
Step 1 (DP on $[N]\setminus S$): For $[N]\setminus S$, each parameter $W_{j}+\text{noise}_{j}$ is $\epsilon_{j}$-DP. By composition, releasing all these parameters is $\sum_{j\in[N]\setminus S}\epsilon_{j}$-DP [42]. Formally, for any adjacent $D,D^{\prime}$ and event $O$:
$\frac{\Pr[\mathcal{M}_{[N]\setminus S}(D)\in O]}{\Pr[\mathcal{M}_{[N]\setminus S}(D^{\prime})\in O]}\leq e^{\sum_{j\in[N]\setminus S}\epsilon_{j}}.$
[42].
Step 2 (Selective Encryption on $S$): Parameters in $S$ are encrypted with a semantically secure HE scheme. Without the secret key, no information about the plaintext is revealed, contributing 0-DP cost [16, 44].
Step 3 (Scrambling as Post-Processing): The scrambling $\mathcal{T}_{k}$ is a deterministic, data-independent post-processing step. A fundamental property of DP is closure under post-processing: if $\mathcal{M}$ is $\epsilon$-DP, then $f(\mathcal{M})$ is also $\epsilon$-DP for any deterministic $f$ [43, 42]. Since $\mathcal{T}_{k}$ does not depend on $D$ beyond $\mathcal{M}(D)$ itself, applying $\mathcal{T}_{k}$ cannot increase or decrease the DP budget.
@@
Thus, $\mathcal{M}^{\prime}(D)$ maintains the same DP guarantee as $\mathcal{M}(D)$. ∎
V Evaluation
This section evaluates our federated learning setup using various security techniques, including homomorphic encryption, differential privacy, and our proposed FAS. We assess their impact on test accuracy, communication cost, computational overhead, and privacy preservation.
Each dataset was split into 10 equal partitions, distributed across machines running Flower clients. Each machine trained its local model over 10 rounds, with 20 epochs per round. A validation set (10% of training data) ensured consistency, and final accuracy was evaluated on a separate test set. All models were trained using CPU resources to simulate realistic constraints.
V-A Datasets Used in Experiments
@@
V-A1 CIFAR-10 Dataset
A benchmark for image classification, CIFAR-10 consists of 60,000 32x32 images across 10 classes. It includes 50,000 training and 10,000 test images, commonly used for model performance evaluation [46].
V-A2 Chest X-ray (COVID-19, Pneumonia) Dataset
This dataset contains 6,339 grayscale X-ray images across three classes: benign, COVID-19, and pneumonia (2,313 images per class). These images aid in lung condition assessment [7].
@@ V-A2 Chest X-ray (COVID-19, Pneumonia) Dataset
Figure 4: Examples from the Chest X-ray (Covid, Pneumonia) dataset [20]
V-A3 CT Kidney Dataset
Comprising 12,446 CT scans categorized as benign (5,077), cyst (3,709), stone (1,377), and tumor (2,283), this dataset aids in kidney disease diagnosis [9].
@@ V-A3 CT Kidney Dataset
Figure 5: Examples from the CT Kidney dataset [20]
V-A4 Diabetic Retinopathy Dataset
This dataset comprises 88,645 fundus images, categorized as benign (65,342) or diabetic retinopathy (23,303). Fundus imaging assists in detecting optic nerve abnormalities [6].
@@ V-A4 Diabetic Retinopathy Dataset
Figure 6: Examples from the Diabetic Retinopathy Detection dataset [20]
V-A5 Lung Cancer Histopathological Images
This dataset contains 15,000 histopathological images, equally distributed across benign, squamous cell carcinoma (SCC), and adenocarcinoma (ACC) classes [8].
@@ V-A5 Lung Cancer Histopathological Images
Figure 7: Examples from the Lung Cancer Histopathological Images dataset. [20]
Our evaluation primarily focused on the CT Kidney dataset, chosen for its stable, reliable results. The Chest X-ray and Diabetic Retinopathy datasets provided additional perspectives to assess model generalizability across different medical imaging contexts. The CT Kidney dataset’s robustness enabled prioritization of security technique evaluation over model-specific tuning.
VI Experimental Setup for Federated Learning
This section details the experimental setup for evaluating various cryptographic techniques in a federated learning context on medical image datasets.
VI-A CloudLab Environment
Experiments were conducted on CloudLab [10], a cloud computing testbed for systems research. Our setup utilized 11 physical machines in a shared-nothing cluster, connected via 10 Gbps Ethernet. Each machine was equipped with dual Intel E5-2683 v3 CPUs (14-core, 2.00 GHz), 256 GB RAM, and two 1 TB hard drives, providing substantial computational resources for federated learning tasks.
VI-B Federated Learning Framework: Flower
Flower [11], an open-source federated learning framework, was used for model training and evaluation, supporting TensorFlow [12], PyTorch [13], and MXNet [14]. One CloudLab machine served as the Flower server, while the remaining 10 functioned as Flower clients.
VI-C Federated Learning Setup
Each dataset was divided into 10 equal partitions, distributed across the Flower clients. The centralized federated learning setup comprised 10 clients and a server. Each model was trained over 10 rounds, with 20 epochs per round. A validation set (10% of training data) monitored performance, and final accuracy was evaluated on a test set. All training used CPUs to simulate resource-constrained environments.
VI-D Rationale for Centralized Federated Learning
Decentralized federated learning was not used due to a lack of mutual trust among clients. A centralized setup ensures data privacy, as clients communicate only with the server, avoiding direct data exchange and aligning with confidentiality requirements.
VI-E Deep Learning Models
@@ VI-E3 EfficientNetB0
VI-E4 MobileNet V2
MobileNet V2 employs lightweight convolutions to reduce computational costs while maintaining accuracy, making it suitable for mobile and embedded applications.
VI-F Encryption Configurations
@@
  • Full-Enc: All model weights encrypted using homomorphic encryption.
  • Partly-Enc: Selective encryption applied to a subset of data, with scrambling for security.
Full Overhead represents additional training time due to full encryption, while Partly Overhead captures the impact of partial encryption.
TABLE I: Training Time (Minutes) and Overheads Across Datasets
@@
Diabetic Retinopathy 1140 673
Table I presents training times (minutes) across datasets. Fully encrypted models require significantly more training time, with ResNet-50 showing the highest overhead. In contrast, partly encrypted models reduce overhead substantially while maintaining security.
@@
Figure 8 shows communication costs across encryption levels, revealing higher costs for fully encrypted models, whereas partly encrypted models optimize performance.
@@
Figure 9 indicates that accuracy remains stable across encryption percentages, confirming the effectiveness of partly encryption in preserving model performance.
Figure 10: Training Time and Encrypted Size Comparison for Fully and Partly Encrypted Models (CT Kidney dataset)
Figure 10 highlights significant reductions in encryption overhead when using partly encryption, making FAS a viable option for resource-limited federated learning scenarios.
VI-G Effects on Security
Before conducting our main security experiments, we implemented several security tests to evaluate the robustness of our encryption methods. These tests included integrity checks, timing attack tests, noise and precision evaluations, and simulations of chosen-ciphertext and chosen-plaintext attacks. Additional differential attack tests and statistical distribution checks were performed to verify that encryption preserved data integrity and privacy. Our findings indicate that 10% partly encryption successfully passed the same security tests as 100% full encryption, demonstrating strong security with reduced computational overhead.
TABLE II: Full vs Partly Encryption: Security Test Results
Test Partly Enc. Full Enc.

### v4

@@ 4.6.1 Homomorphic Encryption
In the context of FL, HE enables secure aggregation of model updates from distributed clients. The mathematical foundation of HE ensures that:
@@
where Enc and Dec refer to the encryption and decryption functions, respectively, and Eval denotes a computation performed on ciphertexts. This capability is essential for privacy-preserving scenarios, as it ensures that plaintext data remains inaccessible to potential adversaries. gentry2009 .
4.6.2 Differential Privacy
Differential Privacy (DP) is a privacy framework that adds carefully calibrated noise to statistical outputs, ensuring that individual data points have minimal impact on the overall resul.
A mechanism $M$ is said to ensure $\epsilon$-differential privacy if, for all subsets of outputs $O$ and for any pair of datasets $D_{1}$ and $D_{2}$ that differ by only one record, the following holds:
@@ 4.6.2 Differential Privacy
where $\epsilon$ represents the privacy parameter. dwork2006calibrating .
In FL, DP adds noise to model updates or gradients prior to aggregation, shielding individual data contributions while retaining overall model performance. This makes DP particularly useful for safeguarding privacy in high-stakes sectors like medical and financial applications.
4.7 Process Flow of the Proposed Model
The proposed model follows a three-stage pipeline: client-side processing, server-side processing, and client post-processing. Each phase includes well-defined steps illustrated with pseudocode and supported by visual diagrams.
Figure 1: General depiction of the FL model showing the processes of encryption, scrambling, and aggregation.
Figure 1 provides a general depiction of our FL model, demonstrating the encryption, scrambling, and noise addition steps across client nodes, and the subsequent aggregation performed by the server.
Figure 2: Client-side process in FL system: encryption, scrambling, and noise addition.
Figure 3: Server-side process in FL system: aggregation and re-scrambling.
Figure 1 presents a general overview of the FL system, encompassing client-side encryption, scrambling, noise addition, and server-side aggregation. Figures 2 and 3 illustrate the detailed client and server sides workflows, respectively.
The process begins at the client side, where local data is utilized to train a model, producing a weight vector. This weight vector is split, encrypted, and further obfuscated through bitwise scrambling and the addition of noise.
Once encrypted, the weight vectors are formatted to hexadecimal and transmitted to the server. The server subsequently collects encrypted model data from all clients and applies federated averaging (FedAvg) to combine the parameters. Importantly, the aggregation is done on both encrypted and non-encrypted vectors, ensuring privacy is maintained while also allowing for more efficient computation.
@@ 4.7 Process Flow of the Proposed Model
This privacy-aware and secure approach facilitates effective model training on decentralized data, while safeguarding sensitive information from unauthorized access.
4.7.1 Client-Side Processing
In the client-side processing stage, model weights ($w$) are prepared for secure transmission to the server by encrypting selected weights, obfuscating the remaining weights, and combining them into a single data structure. Initially, the weights are split into two subsets based on a predefined encryption percentage ($enc\_pct$) (Lines 6–8). The selected portion of the weights is encrypted using a HE function ($enc\_func$), and these encrypted weights are stored in $enc\_w$ (Lines 9–11). For the remaining weights, they are first formatted to resemble encrypted data ($fmt\_weight$), noise is added using $noise\_func$, and the noisy weights are scrambled with a cryptographic key ($scr\_key$) via a scrambling function ($scr\_func$) (Lines 12–16). These scrambled weights are stored in $scr\_w$. Finally, the encrypted and scrambled weights are combined into a single structure ($enc\_scr\_w$) (Line 17) and transmitted to the server (Line 18).
@@ 4.7.1 Client-Side Processing
4.7.2 Server-Side Processing
@@
4.7.3 Client Post-Processing
In the final stage, the client receives the processed weights ($sec\_w$) from the server (Line 2) and restores them for updating the local model. The encrypted subset of weights is decrypted using the appropriate decryption function (Lines 3–4), while the scrambled weights are unscrambled with the cryptographic key ($scr\_key$) to restore their original form (Lines 5–7). These two subsets of weights are then combined to reconstruct the full model weights (Line 8), which are subsequently used to refine the model instance residing on the client for the next round of federated learning. (Line 9).
@@ 4.7.3 Client Post-Processing
15:     Proceed to next round of FL with $final\_w$
@@
4.8 Advantages of the Proposed Model
@@
  • Scalability: Optimized for large-scale FL systems with multiple clients.
4.9 Privacy Analysis
This section presents a provide a formal analysis demonstrating that our FL mechanism—integrating selective homomorphic encryption, differential privacy, and bitwise scrambling—upholds a clear and measurable differential privacy guarantee.
4.9.1 Setup and Definitions
We consider a FL setting with $n$ clients, each holding private data. The global model is described by parameters indexed by $[N]$. Let $S\subseteq[N]$ be the subset of parameters to be protected by selective homomorphic encryption , and let $[N]\setminus S$ be the remaining parameters to which we add differential privacy (DP) noise.
@@ 4.9.1 Setup and Definitions
A mechanism $M$ satisfies $\epsilon$-differential privacy if, for any neighboring datasets $D$ and $D^{\prime}$ (differing in one element), and for every measurable set of outputs $O$, it holds that:
@@
Parameters in $S$ are encrypted using a semantically secure HE scheme. Semantic security ensures no polynomial-time adversary can distinguish ciphertexts of different messages, implying no additional privacy cost in a DP sense (0-DP).gentry2009
@@
For each $i\in[N]\setminus S$, we add noise calibrated to $\epsilon_{i}$-DP. By the composition property of DP, releasing all these parameters together is $\left(\sum_{i\in[N]\setminus S}\epsilon_{i}\right)$-DP.dwork2006calibrating
@@
We define a scrambling function $\mathcal{T}_{k}:\mathcal{X}\to\mathcal{X}$, where $\mathcal{X}$ is the space of possible model outputs. $\mathcal{T}_{k}$ is deterministic, keyed by $k$, and independent of $D$ except through $M(D)$. This scrambling is a form of post-processing.Dwork2006
4.9.2 Privacy Guarantee
Theorem 4.1
@@ 4.9.2 Privacy Guarantee
    Encrypts $W_{i}$ for $i\in S$ using a semantically secure HE scheme gentry2009 ; Fontaine2007 .
@@
    Adds noise to each $W_{j}$ for $j\in[N]\setminus S$ to achieve $\epsilon_{j}$-DP. Releasing all noisy parameters together is $\left(\sum_{j\in[N]\setminus S}\epsilon_{j}\right)$-DP Dwork2014 .
@@
    Applies a scrambling function $\mathcal{T}_{k}$ to the entire output, resulting in $\mathcal{M}^{\prime}(D)=\mathcal{T}_{k}(\mathcal{M}(D))$ Dwork2006 .
@@
Proof
Step 1 (DP on $[N]\setminus S$): For $[N]\setminus S$, each parameter $W_{j}+\text{noise}_{j}$ is $\epsilon_{j}$-DP. By composition, releasing all these parameters is $\sum_{j\in[N]\setminus S}\epsilon_{j}$-DP Dwork2014 . Formally, for any adjacent $D,D^{\prime}$ and event $O$:
[…]
Step 2 (Selective Encryption on $S$): Parameters in $S$ are encrypted with a semantically secure HE scheme. Without the secret key, no information about the plaintext is revealed, contributing 0-DP cost gentry2009 ; Fontaine2007 .
Step 3 (Scrambling as Post-Processing): The scrambling function $\mathcal{T}_{k}$ operates as a deterministic transformation that does not depend on the underlying data. A core principle of differential privacy is that its guarantees are preserved under any data-independent transformation. Specifically, if a mechanism $\mathcal{M}$ satisfies $\epsilon$-DP, then so does $f(\mathcal{M})$ for any deterministic function $f$ Dwork2006 ; Dwork2014 .
@@
Thus, $\mathcal{M}^{\prime}(D)$ maintains the same DP guarantee as $\mathcal{M}(D)$.
5 Evaluation
This section evaluates our FL setup using various security techniques, including HE, differential privacy, and our proposed FAS. We assess their impact on test accuracy, communication cost, computational overhead, and privacy preservation.
Each dataset was divided into 10 equal parts and assigned to separate machines operating as Flower flower clients. Each client trained its local model over 10 rounds, performing 20 epochs per round. A validation set, comprising 10% of the training data, was used to maintain consistency, while final accuracy was measured on an independent test set. All models were trained using CPU resources to simulate realistic constraints.
5.1 Datasets Used in Experiments
@@
5.1.1 CIFAR-10 Dataset CIFAR10
A benchmark for image classification, CIFAR-10 consists of 60,000 32x32 images across 10 classes. It includes 50,000 training and 10,000 test images, commonly used for model performance evaluation.
5.1.2 Chest X-ray (COVID-19, Pneumonia) Dataset ChestXray
This dataset contains 6,339 grayscale X-ray images across three classes: benign, COVID-19, and pneumonia (2,313 images per class). These images aid in lung condition assessment.
@@ 5.1.2 Chest X-ray (COVID-19, Pneumonia) Dataset ChestXray
Figure 4: Examples from the Chest X-ray (Covid, Pneumonia) dataset mine
5.1.3 CT Kidney Dataset kidney
Comprising 12,446 CT scans categorized as benign (5,077), cyst (3,709), stone (1,377), and tumor (2,283), this dataset aids in kidney disease diagnosis.
@@ 5.1.3 CT Kidney Dataset kidney
Figure 5: Examples from the CT Kidney dataset mine
5.1.4 Diabetic Retinopathy Dataset diabetic
This dataset comprises 88,645 fundus images, categorized as benign (65,342) or diabetic retinopathy (23,303). Fundus imaging assists in detecting optic nerve abnormalities.
@@ 5.1.4 Diabetic Retinopathy Dataset diabetic
Figure 6: Examples from the Diabetic Retinopathy Detection dataset mine
5.1.5 Lung Cancer Histopathological Images lungcolon
This dataset contains 15,000 histopathological images, equally distributed across benign, squamous cell carcinoma (SCC), and adenocarcinoma (ACC) classes.
@@ 5.1.5 Lung Cancer Histopathological Images lungcolon
Figure 7: Examples from the Lung Cancer Histopathological Images dataset. mine
6 Experimental Setup
This section outlines the experimental framework employed to assess different cryptographic methods within a FL environment applied to medical imaging datasets.
6.1 CloudLab Environment
Experiments were conducted on CloudLab cloudlab , a cloud computing testbed for systems research. The experimental setup consisted of 11 standalone physical machines in a shared-nothing cluster, interconnected through 10 Gbps Ethernet. Each machine featured dual Intel E5-2683 v3 CPUs (14 cores at 2.00 GHz), 256 GB of RAM, and a pair of 1 TB hard drives, offering ample computational capacity for FE workloads.
6.2 Federated Learning Framework: Flower
Flower flower , an open-source FL framework, was used for model training and evaluation, supporting TensorFlow tensorflow , PyTorch pytorch , and MXNet mxnet . One CloudLab machine served as the Flower server, while the remaining 10 functioned as Flower clients.
6.3 Federated Learning Setup
Each dataset was divided into 10 equal partitions, distributed across the Flower clients. The centralized FL setup comprised 10 clients and a server. Each model was trained over 10 rounds, with 20 epochs per round. A validation set (10% of training data) monitored performance, and final accuracy was evaluated on a test set. All training used CPUs to simulate resource-constrained environments.
6.4 Rationale for Centralized Federated Learning
Decentralized FL was not used due to a lack of mutual trust among clients. A centralized setup ensures data privacy, as clients communicate only with the server, avoiding direct data exchange and aligning with confidentiality requirements.
6.5 Deep Learning Models
@@ 6.5.3 EfficientNetB0efficientnet
6.5.4 MobileNet V2mobilenet
MobileNet V2 leverages efficient convolutional operations to minimize computational overhead while still delivering strong predictive performance, making it well-suited for deployment on mobile and embedded devices.
6.6 Encryption Configurations
@@
  • Full-Enc: All model weights encrypted using HE.
  • Partly-Enc:FAS is applied to a subset of data, with scrambling for security.
Full Overhead represents additional training time due to full encryption, while Partly Overhead captures the impact of FAS.
Table 1: Training Time (Minutes) and Overheads Across Datasets
@@
Diabetic
Retinopathy
1140 673
Table 1 presents training times (minutes) across datasets. Fully encrypted models require significantly more training time, with ResNet-50 showing the highest overhead. In contrast, FAS reduce overhead substantially while maintaining security.
@@
Figure 8 shows communication costs across encryption levels, revealing higher costs for fully encrypted models, whereas FAS optimizes performance.
@@
Figure 9 indicates that accuracy remains stable across encryption percentages, confirming the effectiveness of FAS in preserving model performance.
Figure 10: Training Time and Encrypted Size Comparison for FHE and FAS Models (CT Kidney dataset)
Figure 10 highlights significant reductions in encryption overhead when using FAS, making FAS a viable option for resource-limited FL scenarios.
6.7 Effects on Security
Before conducting our main security experiments, we implemented several security tests to evaluate the robustness of our encryption methods. These tests included integrity checks, timing attack tests, noise and precision evaluations, and simulations of chosen-ciphertext and chosen-plaintext attacks. Additional differential attack tests and statistical distribution checks were performed to verify that encryption preserved data integrity and privacy. Our findings indicate that 10% FAS encryption successfully passed the same security tests as 100% FHE, demonstrating strong security with reduced computational overhead.
Table 2: FHE vs FAS: Security Test Results
Test FAS FHE